\def\soft#1{\leavevmode\setbox0=\hbox{h}\dimen7=\ht0\advance \dimen7
  by-1ex\relax\if t#1\relax\rlap{\raise.6\dimen7
  \hbox{\kern.3ex\char'47}}#1\relax\else\if T#1\relax
  \rlap{\raise.5\dimen7\hbox{\kern1.3ex\char'47}}#1\relax \else\if
  d#1\relax\rlap{\raise.5\dimen7\hbox{\kern.9ex \char'47}}#1\relax\else\if
  D#1\relax\rlap{\raise.5\dimen7 \hbox{\kern1.4ex\char'47}}#1\relax\else\if
  l#1\relax \rlap{\raise.5\dimen7\hbox{\kern.4ex\char'47}}#1\relax \else\if
  L#1\relax\rlap{\raise.5\dimen7\hbox{\kern.7ex
  \char'47}}#1\relax\else\message{accent \string\soft \space #1 not
  defined!}#1\relax\fi\fi\fi\fi\fi\fi}
\begin{thebibliography}{38}
\providecommand{\natexlab}[1]{#1}
\providecommand{\msn}[1]{MR:\,\href{http://www.ams.org/mathscinet-getitem?mr=MR{#1}}{#1}}
\providecommand{\ZBL}[1]{Zbl:\,\href{https://www.zentralblatt-math.org/zmath/en/search/?q=an:#1}{#1}}
\providecommand{\url}[1]{\texttt{#1}}
\providecommand{\urlprefix}{}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi:\discretionary{}{}{}#1}\else
  \providecommand{\doi}{doi:\discretionary{}{}{}\begingroup
  \urlstyle{rm}\Url}\fi

\bibitem[{Alon et~al.(2003)Alon, Ding, Oporowski, and Vertigan}]{ADOV03}
\textsc{Noga Alon, Guoli Ding, Bogdan Oporowski, and Dirk Vertigan}.
\newblock \href{http://dx.doi.org/10.1016/S0095-8956(02)00006-0}{Partitioning
  into graphs with only small components}.
\newblock \emph{J. Combin. Theory Ser. B}, 87(2):231--243, 2003.

\bibitem[{Alon et~al.(1990)Alon, Seymour, and Thomas}]{AST90}
\textsc{Noga Alon, Paul Seymour, and Robin Thomas}.
\newblock \href{http://dx.doi.org/http://dx.doi.org/10.2307/1990903}{A
  separator theorem for nonplanar graphs}.
\newblock \emph{J. Amer. Math. Soc.}, 3(4):801--808, 1990.

\bibitem[{Bar{\'a}t and Wood(2008)}]{BW08}
\textsc{J{\'a}nos Bar{\'a}t and David~R. Wood}.
\newblock \href{http://dx.doi.org/10.37236/823}{Notes on nonrepetitive graph
  colouring}.
\newblock \emph{Electron. J. Combin.}, 15:R99, 2008.

\bibitem[{Bodlaender(1988)}]{Bodlaender-IPL88}
\textsc{Hans~L. Bodlaender}.
\newblock \href{http://dx.doi.org/10.1016/0020-0190(88)90051-8}{The complexity
  of finding uniform emulations on fixed graphs}.
\newblock \emph{Inform. Process. Lett.}, 29(3):137--141, 1988.

\bibitem[{Bodlaender(1990)}]{Bodlaender-IC90}
\textsc{Hans~L. Bodlaender}.
\newblock \href{http://dx.doi.org/10.1016/0890-5401(90)90027-F}{The complexity
  of finding uniform emulations on paths and ring networks}.
\newblock \emph{Inform. and Comput.}, 86(1):87--106, 1990.

\bibitem[{Bodlaender(1998)}]{Bodlaender98}
\textsc{Hans~L. Bodlaender}.
\newblock \href{http://dx.doi.org/10.1016/S0304-3975(97)00228-4}{A partial
  $k$-arboretum of graphs with bounded treewidth}.
\newblock \emph{Theoret. Comput. Sci.}, 209(1-2):1--45, 1998.

\bibitem[{Bodlaender(1999)}]{Bodlaender-DMTCS99}
\textsc{Hans~L. Bodlaender}.
\newblock \href{https://dmtcs.episciences.org/256}{A note on domino treewidth}.
\newblock \emph{Discrete Math. Theor. Comput. Sci.}, 3(4):141--150, 1999.

\bibitem[{Bodlaender and Engelfriet(1997)}]{BodEng-JAlg97}
\textsc{Hans~L. Bodlaender and Joost Engelfriet}.
\newblock \href{http://dx.doi.org/10.1006/jagm.1996.0854}{Domino treewidth}.
\newblock \emph{J. Algorithms}, 24(1):94--123, 1997.

\bibitem[{Bodlaender et~al.(2022)Bodlaender, Groenland, and Jacob}]{BGJ22}
\textsc{Hans~L. Bodlaender, Carla Groenland, and Hugo Jacob}.
\newblock \href{https://doi.org/10.4230/LIPIcs.IPEC.2022.7}{{On the
  Parameterized Complexity of Computing Tree-Partitions}}.
\newblock In \textsc{Holger Dell and Jesper Nederlof}, eds., \emph{Proc. 17th
  International Symposium on Parameterized and Exact Computation (IPEC 2022)},
  vol. 249 of \emph{Leibniz Int'l Proc. Informatics (LIPIcs)}, pp. 7:1--7:20.
  Schloss Dagstuhl, 2022.

\bibitem[{Bodlaender and van Leeuwen(1986)}]{BvL-IC86}
\textsc{Hans~L. Bodlaender and Jan van Leeuwen}.
\newblock \href{http://dx.doi.org/10.1016/S0019-9958(86)80008-0}{Simulation of
  large networks on smaller networks}.
\newblock \emph{Inform. and Control}, 71(3):143--180, 1986.

\bibitem[{Campbell et~al.(2022{\natexlab{a}})Campbell, Clinch, Distel, Gollin,
  Hendrey, Hickingbotham, Huynh, Illingworth, Tamitegama, Tan, and Wood}]{UTW}
\textsc{Rutger Campbell, Katie Clinch, Marc Distel, J.~Pascal Gollin, Kevin
  Hendrey, Robert Hickingbotham, Tony Huynh, Freddie Illingworth, Youri
  Tamitegama, Jane Tan, and David~R. Wood}.
\newblock \href{https://arxiv.org/abs/2206.02395}{Product structure of graph
  classes with bounded treewidth}.
\newblock 2022{\natexlab{a}}, arXiv:2206.02395.

\bibitem[{Campbell et~al.(2022{\natexlab{b}})Campbell, Distel, Gollin, Harvey,
  Hendrey, Hickingbotham, Mohar, and Wood}]{CDGHHHMW}
\textsc{Rutger Campbell, Marc Distel, J.~Pascal Gollin, Daniel~J. Harvey, Kevin
  Hendrey, Robert Hickingbotham, Bojan Mohar, and David~R. Wood}.
\newblock \href{https://arxiv.org/abs/2210.13720}{Graphs of linear growth have
  bounded treewidth}.
\newblock 2022{\natexlab{b}}.

\bibitem[{Carmi et~al.(2008)Carmi, Dujmovi\'c, Morin, and Wood}]{CDMW08}
\textsc{Paz Carmi, Vida Dujmovi\'c, Pat Morin, and David~R. Wood}.
\newblock \href{https://www.combinatorics.org/v15i1r107}{Distinct distances in
  graph drawings}.
\newblock \emph{Electron. J. Combin.}, 15:R107, 2008.

\bibitem[{Chatzidimitriou et~al.(2018)Chatzidimitriou, Raymond, Sau, and
  Thilikos}]{CRST18}
\textsc{Dimitris Chatzidimitriou, Jean{-}Florent Raymond, Ignasi Sau, and
  Dimitrios~M. Thilikos}.
\newblock \href{https://doi.org/10.1007/s00453-017-0313-5}{An {O(log
  OPT)}-approximation for covering and packing minor models of $\theta_r$}.
\newblock \emph{Algorithmica}, 80(4):1330--1356, 2018.

\bibitem[{Di~Giacomo et~al.(2005)Di~Giacomo, Liotta, and Meijer}]{GLM05}
\textsc{Emilio Di~Giacomo, Giuseppe Liotta, and Henk Meijer}.
\newblock \href{http://dx.doi.org/10.1016/j.comgeo.2004.11.003}{Computing
  straight-line {3D} grid drawings of graphs in linear volume}.
\newblock \emph{Comput. Geom. Theory Appl.}, 32(1):26--58, 2005.

\bibitem[{Ding and Oporowski(1995)}]{DO95}
\textsc{Guoli Ding and Bogdan Oporowski}.
\newblock \href{http://dx.doi.org/10.1002/jgt.3190200412}{Some results on tree
  decomposition of graphs}.
\newblock \emph{J. Graph Theory}, 20(4):481--499, 1995.

\bibitem[{Ding and Oporowski(1996)}]{DO96}
\textsc{Guoli Ding and Bogdan Oporowski}.
\newblock \href{http://dx.doi.org/10.1016/0012-365X(94)00337-I}{On
  tree-partitions of graphs}.
\newblock \emph{Discrete Math.}, 149(1--3):45--58, 1996.

\bibitem[{Draganić et~al.(2023)Draganić, Kaufmann, Correia, Petrova, and
  Steiner}]{DKCPS}
\textsc{Nemanja Draganić, Marc Kaufmann, David~Munhá Correia, Kalina Petrova,
  and Raphael Steiner}.
\newblock \href{http://arxiv.org/abs/2307.12028}{Size-{R}amsey numbers of
  structurally sparse graphs}.
\newblock 2023, arXiv:2307.12028.

\bibitem[{Dujmovi{\'c} et~al.(2022)Dujmovi{\'c}, Esperet, Morin, Walczak, and
  Wood}]{DEMWW22}
\textsc{Vida Dujmovi{\'c}, Louis Esperet, Pat Morin, Bartosz Walczak, and
  David~R. Wood}.
\newblock \href{http://dx.doi.org/10.1017/S0963548321000213}{Clustered
  3-colouring graphs of bounded degree}.
\newblock \emph{Combin. Probab. Comput.}, 31(1):123--135, 2022.

\bibitem[{Dujmovi\'c et~al.(2022)Dujmovi\'c, Joret, Micek, Morin, and
  Wood}]{DJMMW}
\textsc{Vida Dujmovi\'c, Gwena\"el Joret, Piotr Micek, Pat Morin, and David~R.
  Wood}.
\newblock \href{https://arxiv.org/abs/2212.02388}{Bad news for product
  structure of bounded-degree graphs}.
\newblock 2022.

\bibitem[{Dujmovi{\'c} et~al.(2005)Dujmovi{\'c}, Morin, and Wood}]{DMW05}
\textsc{Vida Dujmovi{\'c}, Pat Morin, and David~R. Wood}.
\newblock \href{http://dx.doi.org/10.1137/S0097539702416141}{Layout of graphs
  with bounded tree-width}.
\newblock \emph{SIAM J. Comput.}, 34(3):553--579, 2005.

\bibitem[{Dujmovi{\'c} et~al.(2007)Dujmovi{\'c}, Suderman, and Wood}]{DSW07}
\textsc{Vida Dujmovi{\'c}, Matthew Suderman, and David~R. Wood}.
\newblock \href{http://dx.doi.org/10.1016/j.comgeo.2006.08.002}{Graph drawings
  with few slopes}.
\newblock \emph{Comput. Geom. Theory Appl.}, 38:181--193, 2007.

\bibitem[{Edenbrandt(1986)}]{Edenbrandt86}
\textsc{Anders Edenbrandt}.
\newblock \href{https://doi.org/http://dx.doi.org/10.1007/BF01933740}{Quotient
  tree partitioning of undirected graphs}.
\newblock \emph{BIT}, 26(2):148--155, 1986.

\bibitem[{Fishburn and Finkel(1982)}]{FF82}
\textsc{John~P. Fishburn and Raphael~A. Finkel}.
\newblock \href{http://dx.doi.org/10.1109/TC.1982.1675994}{Quotient networks}.
\newblock \emph{IEEE Trans. Comput.}, C-31(4):288--295, 1982.

\bibitem[{Giannopoulou et~al.(2016)Giannopoulou, Kwon, Raymond, and
  Thilikos}]{GKRT16}
\textsc{Archontia~C. Giannopoulou, O{-}joung Kwon, Jean{-}Florent Raymond, and
  Dimitrios~M. Thilikos}.
\newblock \href{https://doi.org/10.1007/978-3-662-53536-3\_7}{Packing and
  covering immersion models of planar subcubic graphs}.
\newblock In \textsc{Pinar Heggernes}, ed., \emph{Proc. 42nd Int'l Workshop on
  Graph-Theoretic Concepts in Computer Science \textup{({WG} 2016)}}, vol. 9941
  of \emph{Lecture Notes in Comput. Sci.}, pp. 74--84. 2016.

\bibitem[{Halin(1991)}]{Halin91}
\textsc{Rudolf Halin}.
\newblock \href{http://dx.doi.org/10.1016/0012-365X(91)90436-6}{Tree-partitions
  of infinite graphs}.
\newblock \emph{Discrete Math.}, 97:203--217, 1991.

\bibitem[{Harvey and Wood(2017)}]{HW17}
\textsc{Daniel~J. Harvey and David~R. Wood}.
\newblock \href{http://dx.doi.org/10.1002/jgt.22030}{Parameters tied to
  treewidth}.
\newblock \emph{J. Graph Theory}, 84(4):364--385, 2017.

\bibitem[{Kamcev et~al.(2021)Kamcev, Liebenau, Wood, and Yepremyan}]{KLWY21}
\textsc{Nina Kamcev, Anita Liebenau, David~R. Wood, and Liana Yepremyan}.
\newblock \href{http://dx.doi.org/10.1137/20M1335790}{The size {R}amsey number
  of graphs with bounded treewidth}.
\newblock \emph{SIAM J. Discrete Math.}, 35(1):281--293, 2021.

\bibitem[{Kuske and Lohrey(2005)}]{KuskeLohrey05}
\textsc{Dietrich Kuske and Markus Lohrey}.
\newblock \href{http://dx.doi.org/10.1016/j.apal.2004.06.002}{Logical aspects
  of {C}ayley-graphs: the group case}.
\newblock \emph{Ann. Pure Appl. Logic}, 131(1--3):263--286, 2005.

\bibitem[{Liu and Oum(2018)}]{LO18}
\textsc{Chun-Hung Liu and Sang-il Oum}.
\newblock \href{http://dx.doi.org/10.1016/j.jctb.2017.08.003}{Partitioning
  {$H$}-minor free graphs into three subgraphs with no large components}.
\newblock \emph{J. Combin. Theory Ser. B}, 128:114--133, 2018.

\bibitem[{Raymond and Thilikos(2017)}]{RT17}
\textsc{Jean-Florent Raymond and Dimitrios~M. Thilikos}.
\newblock \href{http://dx.doi.org/10.1016/j.dam.2016.12.025}{Recent techniques
  and results on the {Erd\H{o}s-P{\'{o}}sa} property}.
\newblock \emph{Discrete Appl. Math.}, 231:25--43, 2017.

\bibitem[{Reed(2003)}]{Reed03}
\textsc{Bruce~A. Reed}.
\newblock \href{http://dx.doi.org/10.1007/0-387-22444-0\_4}{Algorithmic aspects
  of tree width}.
\newblock In \emph{Recent advances in algorithms and combinatorics}, vol.~11,
  pp. 85--107. Springer, 2003.

\bibitem[{Robertson and Seymour(1986)}]{RS-II}
\textsc{Neil Robertson and Paul Seymour}.
\newblock \href{https://doi.org/10.1016/0196-6774(86)90023-4}{Graph minors.
  {II}. {A}lgorithmic aspects of tree-width}.
\newblock \emph{J. Algorithms}, 7(3):309--322, 1986.

\bibitem[{Seese(1985)}]{Seese85}
\textsc{Detlef Seese}.
\newblock \href{http://dx.doi.org/10.1007/BFb0028825}{Tree-partite graphs and
  the complexity of algorithms}.
\newblock In \textsc{Lothar Budach}, ed., \emph{Proc. Int'l Conf. on
  Fundamentals of Computation Theory}, vol. 199 of \emph{Lecture Notes in
  Comput. Sci.}, pp. 412--421. Springer, 1985.

\bibitem[{Wood(2006)}]{Wood06}
\textsc{David~R. Wood}.
\newblock \href{https://doi.org/http://dx.doi.org/10.1002/jgt.20183}{Vertex
  partitions of chordal graphs}.
\newblock \emph{J. Graph Theory}, 53(2):167--172, 2006.

\bibitem[{Wood(2009)}]{Wood09}
\textsc{David~R. Wood}.
\newblock \href{http://dx.doi.org/10.1016/j.ejc.2008.11.010}{On
  tree-partition-width}.
\newblock \emph{European J. Combin.}, 30(5):1245--1253, 2009.

\bibitem[{Wood and Telle(2007)}]{WT07}
\textsc{David~R. Wood and Jan~Arne Telle}.
\newblock \href{http://nyjm.albany.edu/j/2007/13-8.html}{Planar decompositions
  and the crossing number of graphs with an excluded minor}.
\newblock \emph{New York J. Math.}, 13:117--146, 2007.

\bibitem[{Zhang and Amini(2022)}]{ZA22}
\textsc{Rui-Ray Zhang and Massih-Reza Amini}.
\newblock \href{https://arxiv.org/abs/2203.13534}{Generalization bounds for
  learning under graph-dependence: A survey}.
\newblock 2022.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage[T1]{fontenc}
\usepackage{iwona,amsmath,amsthm,amsfonts,amssymb,microtype,thmtools,underscore,mathtools,anyfontsize} 
\usepackage[usenames,dvipsnames,svgnames,table]{xcolor}
\usepackage[unicode=true]{hyperref}
\hypersetup{%
colorlinks,
breaklinks=true,
linkcolor={black},
citecolor={black},
urlcolor={blue!60!black},
pdftitle={Tree-Partitions with Small Bounded Degree Trees},
pdfauthor={Distel--Wood}
}
\usepackage[noabbrev,capitalise]{cleveref}
\crefname{lem}{Lemma}{Lemmas}
\crefname{thm}{Theorem}{Theorems}
\crefname{cor}{Corollary}{Corollaries}
\crefname{prop}{Proposition}{Propositions}
\crefname{conj}{Conjecture}{Conjectures}
\crefname{rmk}{Remark}{Remarks}
\crefname{openproblem}{Open Problem}{Open Problems}
\crefformat{equation}{(#2#1#3)}
\Crefformat{equation}{Equation #2(#1)#3}
\usepackage[shortlabels]{enumitem}
\setlist[itemize]{topsep=0ex,itemsep=0ex,parsep=0.4ex}
\setlist[enumerate]{topsep=0ex,itemsep=0ex,parsep=0.4ex}
\usepackage[longnamesfirst,numbers,sort&compress]{natbib}
\makeatletter
\def\NAT@spacechar{~}
\makeatother
\usepackage[margin=30mm]{geometry}
\renewcommand{\baselinestretch}{1.1}
\allowdisplaybreaks
\newcommand{\defn}[1]{\textcolor{Maroon}{\emph{#1}}}

\newcommand{\NN}{\mathbb{N}}

\renewcommand{\geq}{\geqslant}
\renewcommand{\leq}{\leqslant}

\DeclareMathOperator{\tw}{tw}
\DeclareMathOperator{\tpw}{tpw}

\renewcommand{\thefootnote}{\fnsymbol{footnote}}
\allowdisplaybreaks
\theoremstyle{plain}
\newtheorem{thm}{Theorem}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{cor}[thm]{Corollary}

\theoremstyle{definition}

\begin{document}

\title{\bf\Large Tree-Partitions with Small Bounded Degree Trees\footnote{\today. This work was initiated at the \href{https://www.matrix-inst.org.au/events/structural-graph-theory-downunder-ll/}{Structural Graph Theory Downunder II} workshop at the Mathematical Research Institute MATRIX (March 2022). A preliminary version of this paper appears in the 
\href{https://www.matrix-inst.org.au/2021-matrix-annals/}{2021-22 MATRIX Annals}.}}
\author{Marc Distel\,\footnotemark[3] \qquad David~R.~Wood\,\footnotemark[3]}

\footnotetext[3]{School of Mathematics, Monash University, Melbourne, Australia (\texttt{\{marc.distel,david.wood\}@monash.edu}). Research of D.W.\ supported by the Australian Research Council. Research of M.D.\ supported by an Australian Government Research Training Program Scholarship.}

\date{}
\maketitle

\begin{abstract}
A \defn{tree-partition} of a graph $G$ is a partition of $V(G)$ such that identifying the vertices in each part gives a tree. It is known that every graph with treewidth $k$ and maximum degree $\Delta$ has a tree-partition with parts of size $O(k\Delta)$. We prove the same result with the extra property that the underlying tree has maximum degree $O(\Delta)$ and $O(|V(G)|/k)$ vertices. 
\end{abstract}


\renewcommand{\thefootnote}{\arabic{footnote}}

\section{Introduction}

For a graph~$G$ and a tree~$T$, a \defn{$T$-partition} of~$G$ is a partition~${(V_x \colon x\in V(T))}$ of~${V(G)}$ indexed by the nodes of~$T$, such that for every edge~${vw}$ of~$G$, if~${v \in V_x}$ and~${w \in V_y}$, then~${x = y}$ or~${xy \in E(T)}$. The \defn{width} of a $T$-partition is~${\max\{ |{V_x}| \colon x \in V(T)\}}$. The \defn{tree-partition-width}\footnote{Tree-partition-width has also been called \defn{strong treewidth} \citep{BodEng-JAlg97,Seese85}.} of a graph $G$ is the minimum width of a tree-partition of $G$. 

Tree-partitions were independently introduced by \citet{Seese85} and \citet{Halin91}, and have since been widely investigated \citep{Bodlaender-DMTCS99,BodEng-JAlg97, DO95,DO96,Edenbrandt86, Wood06,Wood09,BGJ22}. 
Applications of tree-partitions include 
graph drawing~\citep{CDMW08,GLM05,DMW05,DSW07,WT07}, 
graphs of linear growth~\citep{CDGHHHMW}, 
nonrepetitive graph colouring~\citep{BW08}, 
clustered graph colouring~\citep{ADOV03,LO18}, 
monadic second-order logic~\citep{KuskeLohrey05}, 
network emulations~\citep{Bodlaender-IPL88, Bodlaender-IC90, BvL-IC86, FF82}, statistical learning theory~\citep{ZA22}, 
size Ramsey numbers~\citep{DKCPS,KLWY21}, 
and the edge-{E}rd{\H{o}}s-{P}{\'o}sa property~\citep {RT17,GKRT16,CRST18}. Tree-partitions are also related to graph product structure theory since a graph $G$ has a $T$-partition of width at most $k$ if and only if $G$ is isomorphic to a subgraph of $T\boxtimes K_k$ for some tree $T$; see~\citep{UTW,DJMMW,DEMWW22} for example.

Bounded tree-partition-width implies bounded treewidth\footnote{A \defn{tree-decomposition} of a graph $G$ is a collection $(B_x\subseteq V(G):x\in V(T))$ of subsets of $V(G)$ (called \defn{bags}) indexed by the nodes of a tree $T$, such that: (a) for every edge $uv\in E(G)$, some bag $B_x$ contains both $u$ and $v$; and (b) for every vertex $v\in V(G)$, the set $\{x\in V(T):v\in B_x\}$ induces a non-empty subtree of $T$. The \defn{width} of a tree-decomposition is the size of the largest bag, minus $1$. 
The \defn{treewidth} of a graph $G$, denoted by \defn{$\tw(G)$}, is the minimum width of a tree-decomposition of $G$. Treewidth is the standard measure of how similar a graph is to a tree. Indeed, a connected graph has treewidth 1 if and only if it is a tree. Treewidth is of fundamental importance in structural and algorithmic graph theory; see \citep{Reed03,HW17,Bodlaender98} for surveys.}, as noted by \citet{Seese85}. In particular, for every graph~$G$,
\[\tw(G) \leq 2\tpw(G)-1.\]
Of course, ${\tw(T) = \tpw(T) = 1}$ for every tree~$T$. 
But in general, $\tpw(G)$ can be much larger than~$\tw(G)$. 
For example, fan graphs on~$n$ vertices have treewidth~2 and tree-partition-width~$\Omega(\sqrt{n})$. 
On the other hand, the referee of \citep{DO95} showed that if the maximum degree and treewidth are both bounded, then so is the tree-partition-width, which is one of the most useful results about tree-partitions. A graph $G$ is \defn{trivial} if $E(G)=\emptyset$. Let \defn{$\Delta(G)$} be the maximum degree of $G$. 
 
\begin{thm}[\cite{DO95}]
\label{TreeProduct}
For any non-trivial graph $G$,
$$\tpw(G) \leq 24 (\tw(G)+1)\Delta(G).$$
\end{thm}

\cref{TreeProduct} is stated in \cite{DO95} with ``$\tw(G)$'' instead of ``$\tw(G)+1$'', but a close inspection of the proof shows that ``$\tw(G)+1$'' is needed. \citet{Wood09} showed that \cref{TreeProduct} is best possible up to the multiplicative constant, and also improved the constant $24$ to $9+6\sqrt{2}\approx 17.48$.

This paper proves similar results to \cref{TreeProduct} where the tree $T$ indexing the tree-partition has two extra properties. 

The first property is the maximum degree of $T$. Consider a tree-partition $(B_x:x\in V(T))$ of a graph $G$ with width $k$. For each node $x\in V(T)$, there are at most $\sum_{v\in B_x}\deg(v)$ edges between $B_x$ and $G-B_x$. Thus we may assume that $\deg_T(x)\leq |B_x|\Delta(G)\leq k\Delta(G)$, otherwise delete an `unused' edge of $T$ and add an edge to $T$ between leaf vertices of the resulting component subtrees. It follows that if $\tpw(G)\leq k$ then $G$ has a $T$-partition of width at most $k$ for some tree $T$ with maximum degree at most $\max\{k\Delta(G),2\}$. By \cref{TreeProduct}, every graph $G$ has a $T$-partition of width at most $24 (\tw(G)+1)\Delta(G)$ for some tree $T$ with maximum degree at most $24 (\tw(G)+1)\Delta(G)^2$. This fact has been used in several applications of \cref{TreeProduct} (see \citep{CDMW08,DSW07} for example). 

The second property is the number of vertices in $T$. This property is motivated by results about size-Ramsey number by \citet{DKCPS}, where tree-partitions of $n$-vertex graphs with $\tw(G)\in O(\sqrt{n})$ play a critical role, and it is essential that $\Delta(T)$ is independent of $\tw(G)$ and $|V(T)| \ll |V(G)|$. 

The following theorem improves the above-mentioned upper bound on $\Delta(T)$ so that it is independent of $\tw(G)$, and provides a non-trivial upper bound on $|V(T)|$. Let $\NN=\{1,2,\dots\}$. 

\begin{thm}
\label{ImprovedTreeProduct}
For any integers $k,d\in\NN$, every non-trivial graph $G$ with $\tw(G)\leq k-1$ and $\Delta(G)\leq d$ has a $T$-partition of width at most \(18kd\), for some tree $T$ with $\Delta(T)\leq 6d$ and $|V(T)|\leq\max\{\frac{|V(G)|}{2k},1\}$.
\end{thm}

\cref{ImprovedTreeProduct} enables a $\tw(G)\Delta(G)^2$ term to be replaced by a $\Delta(G)$ term in various  results~\citep{CDMW08,DSW07}. And since $\Delta(T)$ is independent of $\tw(G)$, and $|V(T)|\ll |V(G)|$ when $\tw(G)$ is unbounded, \cref{ImprovedTreeProduct} is suitable for the applications to size-Ramsey numbers in  \citep{DKCPS}. Indeed, \cref{ImprovedTreeProduct} improves upon a similar result of \citet[Lemma~2.1]{DKCPS} which has an extra $O(\log n)$ factor in the width.

Here we give an example of \cref{ImprovedTreeProduct}. \citet{AST90} proved that every $K_t$-minor-free graph $G$ has treewidth at most  $t^{3/2}|V(G)|^{1/2}-1$. \cref{ImprovedTreeProduct} thus implies:

\begin{cor}
\label{MinorClosedClass}
Every $K_t$-minor-free graph with maximum degree $\Delta$ has a $T$-partition of width at most $18t^{3/2}\Delta|V(G)|^{1/2}$, for some tree $T$ with $\Delta(T)\leq 6\Delta$ and $|V(T)|\leq |V(G)|^{1/2}/2t^{3/2}$.
\end{cor}

As mentioned above, \citet{Wood09} improved the constant $24$ to $9+6\sqrt{2}$ in \cref{TreeProduct}. By tweaking the constants in the proof of \cref{ImprovedTreeProduct}, we match this constant with a small increase in the bound on $\Delta(T)$; see \cref{tweaking}.
We choose to first present the proof with integer coefficients for ease of understanding.

Our final result shows that the linear upper bound on $\Delta(T)$ in \cref{ImprovedTreeProduct} is best possible even for trees.

\begin{restatable}{prop}{LowerBound}
\label{LowerBound}
For any integer $\Delta\geq 3$ there exist $\alpha>0$ such that there are infinitely many trees $X$ with maximum degree $\Delta$ such that for every tree $T$ with maximum degree less than $\Delta$, every $T$-partition of $X$ has width at least $|V(X)|^\alpha$. Moreover, if $\Delta=3$ then $\alpha$ can be taken to be arbitrarily close to $1$. 
\end{restatable}

\section{Main Proofs}

The proof of \cref{ImprovedTreeProduct} closely follows the proof of \cref{TreeProduct}, except that we pay attention to $\Delta(T)$ and $|V(T)|$. 
The next lemma is the heart of our proof. 

\begin{lem}
\label{heart}
For $k,d\in\NN$, for any graph $G$ with $\tw(G)\leq k-1$ and $\Delta(G)\leq d$, for any set $S\subseteq V(G)$ with $4 k\leq|S| \leq 12 kd$, there exists a tree-partition $(B_x:x\in V(T))$ of $G$ such that:
\begin{itemize}
    \item $|B_x|\leq 18kd $ for each $x\in V(T)$,
    \item $\Delta(T)\leq 6d$, 
    \item $|V(T)|\leq \frac{|V(G)|}{2k}$.
\end{itemize}
Moreover, there exists $z\in V(T)$ such that:
\begin{itemize}
    \item $S\subseteq B_z$, 
    \item $|B_z|\leq \frac32|S|-2k$,
    \item $\deg_T(z)\leq \frac{|S|}{2k} - 1$.
\end{itemize}
\end{lem}

\begin{proof}
We proceed by induction on $|V(G)$|.

\textbf{Case 1.} $|V(G-S)|\leq 18kd$: Let $T$ be the tree with $V(T)=\{y,z\}$ and $E(T)=\{yz\}$. Note that $\Delta(T)=1\leq 6 d$ and $|V(T)|=2\leq \frac{|S|}{2k} \leq \frac{|V(G)|}{2k}$ and $\deg_T(z)=1\leq \frac{|S|}{2k}-1$. Let $B_z:=S$ and $B_y:=V(G-S)$. Thus $|B_z|=|S|\leq \frac32 |S|-2k\leq 18kd$ and $|B_y|\leq |V(G-S)|\leq 18kd$. Hence $(B_x:x\in V(T))$ is the desired tree-partition of $G$. 

Now assume that $|V(G-S)|\geq 18kd$.

\textbf{Case 2.} $4k \leq |S|\leq 12 k$: Let $S':=\bigcup\{ N_G(v)\setminus S: v\in S\}$. Thus $|S'|\leq d |S|\leq 12kd$. If $|S'|< 4 k$ then add $4k-|S'|$ vertices from $V(G-S-S')$ to $S'$, so that $|S'|=4k$. This is well-defined since 
$|V(G-S)| \geq 18kd \geq 4k$, implying $|V(G-S-S')| \geq 4k-|S'|$.
By induction, there exists a tree-partition $(B_x:x\in V(T'))$ of $G-S$ 
such that:
\begin{itemize}
    \item $|B_x|\leq 18kd $ for each $x\in V(T'))$,
    \item $\Delta(T')\leq 6d$, 
    \item $|V(T')|\leq \frac{|V(G-S)|}{2k}$.
\end{itemize}
Moreover, there exists $z'\in V(T')$ such that:
\begin{itemize}
    \item $S'\subseteq B_{z'}$, 
    \item $|B_{z'}|\leq \frac32 |S'|-2k \leq 18kd -2k$,
    \item $\deg_{T'}(z')\leq \frac{|S'|}{2k} - 1 \leq 6d - 1$.
\end{itemize}
Let $T$ be the tree obtained from $T'$ by adding one new node $z$ adjacent to $z'$. Let $B_z:=S$. So $(B_x:x\in V(T))$ is a tree-partition of $G$ with width at most $\max\{18kd,|S|\}\leq\max\{18kd,12k\}=18kd$. By construction, $\deg_T(z)=1 \leq \frac{|S|}{2k} - 1$ and $\deg_{T}(z') = \deg_{T'}(z')+1\leq (6d-1) + 1 = 6d$. Every other vertex in $T$ has the same degree as in $T'$. Hence $\Delta(T)\leq 6d$, as desired. 
Also $|V(T)|=|V(T')|+1\leq \frac{|V(G-S)|}{2k} +1 <\frac{|V(G)|}{2k}$ since $|S|\geq 4k$. Finally, $S=B_z$ and $|B_z|=|S| \leq \frac32 |S|-2k$.

\textbf{Case 3.} $12 k \leq |S|\leq 12kd$: By the separator lemma of \citet[(2.6)]{RS-II}, there are induced subgraphs $G_1$ and $G_2$ of $G$ with $G_1\cup G_2=G$ and $|V(G_1\cap G_2)|\leq k$, where $|S\cap V(G_i)|\leq \frac23 |S|$ for each $i\in\{1,2\}$. Let $S_i := (S\cap V(G_i))\cup V(G_1\cap G_2)$ for each $i\in\{1,2\}$.

We now bound $|S_i|$. For a lower bound, since $|S\cap V(G_1)|\leq \frac23 |S|$, we have $|S_2|\geq |S\setminus V(G_1)|\geq \frac13 |S| \geq 4k $. By symmetry, $|S_1|\geq  4k $. For an upper bound, $|S_i|\leq\frac23 |S| + k \leq 8kd + k \leq 12kd$. Also note that $|S_1|+|S_2|\leq |S|+2k$.

We have shown that $4k \leq |S_i|\leq 12kd$ for each $i\in\{1,2\}$. Thus we may apply induction to $G_i$ with $S_i$ the specified set. Hence there exists a tree-partition $(B^i_x:x\in V(T_i))$ of $G_i$ 
such that:
\begin{itemize}
    \item $|B^i_x|\leq 18kd $ for each $x\in V(T_i))$,
    \item $\Delta(T_i)\leq 6d$, 
    \item $|V(T_i)|\leq \frac{|V(G_i)|}{2k}$.
\end{itemize}
Moreover, there exists $z_i\in V(T_i)$ such that:
\begin{itemize}
    \item $S_i\subseteq B_{z_i}$, 
    \item $|B_{z_i}|\leq \frac32|S_i|-2k$,
    \item $\deg_{T_i}(z_i)\leq \frac{|S_i|}{2k}-1$.
\end{itemize}
Let $T$ be the tree obtained from the disjoint union of $T_1$ and $T_2$ by identifying $z_1$ and $z_2$ into a vertex $z$. Let $B_z:= B^1_{z_1}\cup B^2_{z_2}$. Let $B_x:= B^i_x$ for each $x\in V(T_i)\setminus\{z_i\}$. Since $G=G_1\cup G_2$ and $V(G_1\cap G_2)\subseteq B^1_{z_1}\cap B^2_{z_2} \subseteq B_z$, we have that $(B_x:x\in V(T))$ is a tree-partition of $G$. 
By construction, $S\subseteq B_z$ and since $V(G_1\cap G_2)\subseteq B^i_{z_i}$ for each $i$, 
\begin{align*}
    |B_z| 
    & \leq |B^1_{z_1}|+|B^2_{z_2}| - |V(G_1\cap G_2)|\\
    & \leq (\tfrac32|S_1|-2k) +  (\tfrac32|S_2|-2k) - |V(G_1\cap G_2)|\\
    & = \tfrac32( |S_1|+ |S_2|) -4k - |V(G_1\cap G_2)|\\
    & \leq \tfrac32( |S| + 2|V(G_1\cap G_2)| ) -4k - |V(G_1\cap G_2)|\\
    & \leq \tfrac32 |S| + 2|V(G_1\cap G_2)| -4k\\
    & \leq \tfrac32 |S| - 2 k\\
    & < 18 kd.
\end{align*}
Every other part has the same size as in the tree-partition of $G_1$ or $G_2$. So this tree-partition of $G$ has width at most $18kd$. 
Note that 
\begin{align*}
 \deg_T(z)   = \deg_{T_1}(z_1) + \deg_{T_2}(z_2)
    & \leq  (\tfrac{|S_1|}{2k}-1) + (\tfrac{|S_2|}{2k}-1)\\
    & =  \tfrac{|S_1|+|S_2|}{2k} -2\\
    & \leq  \tfrac{|S|+2k}{2k} -2\\
    & =  \tfrac{|S|}{2k} -1\\
    & < 6 d. 
\end{align*}
Every other node of $T$ has the same degree as in $T_1$ or $T_2$. 
Thus $\Delta(T) \leq 6d$. 
Finally, 
\begin{align*}
    |V(T)|  =|V(T_1)|+|V(T_2)|-1 
    & \leq \tfrac{|V(G_1)|}{2k}+\tfrac{|V(G_2)|}{2k}-1\\
    & \leq  \tfrac{|V(G)|+k}{2k}-1\\
    & < \tfrac{|V(G)|}{2k}.
\end{align*}
This completes the proof.
\end{proof}

We now prove our main result. 

\begin{proof}[Proof of \cref{ImprovedTreeProduct}]
First suppose that $|V(G)| < 4 k$. Let $T$ be the 1-vertex tree with $V(T)=\{x\}$, and let $B_x:=V(G)$. Then $(B_x:x\in V(T))$ is the desired tree-partition, since $|V(T)|= 1 \leq \max\{\frac{|V(G)|}{2k},1\}$ and $|B_x|=|V(G)|<4 k \leq 18 kd$ and $\Delta(T)=0\leq 6d$. 
Now assume that $|V(G)| \geq 4 k$. The result follows from \cref{heart}, where $S$ is any set of $4k$ vertices in $G$. 
\end{proof}

We now prove the lower bound. For $\Delta,d\in\NN$ with $\Delta\geq 2$, let $X_{\Delta,d}$ be the tree rooted at a vertex $r$ such that every leaf is at distance $d$ from $r$ and every non-leaf vertex has degree $\Delta$. Observe that $X_{\Delta,d}$ has the maximum number of vertices in a tree with maximum degree $\Delta$ and radius $d$, where  
\[ |V(X_{\Delta,d})| 
= 1+\Delta\sum_{i=0}^{d-1}(\Delta-1)^i.\]
Note that $|V(X_{2,d})|= 2d+1$, and
 if $\Delta\geq 3$ then 
\[ (\Delta-1)^d \leq |V(X_{\Delta,d})| 
=1+\tfrac{\Delta}{\Delta-2}\big((\Delta-1)^d-1\big)
\leq 3(\Delta-1)^d.\]

\LowerBound*

\begin{proof}
First suppose that $\Delta\geq 4$. 
Let $d_0\in\NN$ be sufficiently large so that $(\frac{\Delta-1}{\Delta-2})^{d_0}>3$. 
Let $\alpha:= {1- \log_{\Delta-1}(3^{1/d_0}(\Delta-2))}$, which is positive by the choice of $d_0$. 
Let $d\in\NN$ with $d\geq d_0$. 
It follows that 
$(\Delta-1)^{(1-\alpha)d} \geq 3(\Delta-2)^d $. 
Consider any tree-partition $(B_u:u\in V(T))$ of $X_{\Delta,d}$, where $T$ is any tree with maximum degree at most $\Delta-1$. Let $z$ be the vertex of $T$ such that the root $r\in B_z$. Since adjacent vertices in $X_{\Delta,d}$ belong to adjacent parts or the same part in $T$, every vertex in $T$ is at distance at most $d$ from $z$. Thus $T$ has radius at most $d$, and \[|V(T)|\leq|V(X_{\Delta-1,d})| 
  \leq 3(\Delta-2)^d 
  \leq (\Delta-1)^{(1-\alpha)d} \leq
|V(X_{\Delta,d})|^{1-\alpha}  .\]
By the pigeon-hole principle, there is a vertex $u\in V(T)$ such that 
$|B_u|\geq \frac{|V(X_{\Delta,d})|}{|V(T)|}\geq |V(X_{\Delta,d})|^\alpha$.

Now assume that $\Delta=3$. Let $\alpha\in(0,1)$, let $d_0\in\NN$ be sufficiently large so that $2d_0+1 \leq 2^{(1-\alpha)d_0}$, and let $d\geq d_0$. So $2d+1 \leq 2^{(1-\alpha)d}$. Consider any tree-partition $(B_u:u\in V(T))$ of $X_{3,d}$, where $T$ is any tree with maximum degree at most $2$. By the argument above, $T$ has radius at most $d$, implying \[|V(T)|\leq|V(X_{2,d})|=2d+1 \leq 2^{(1-\alpha)d} \leq |V(X_{3,d})|^{1-\alpha}.\]
Again, there is a vertex $u\in V(T)$ such that 
$|B_u|\geq \frac{|V(X_{3,d})|}{|V(T)|}\geq |V(X_{3,d})|^{\alpha}$.
\end{proof}

\section{Tweaking the Constants} 
\label{tweaking}

Consider the following generalisation of \cref{heart}.

\begin{lem}
\label{alpha}
Fix $k,d\in\NN$ and $\alpha\in\mathbb{R}$ with $\alpha>2$. 
For any  graph $G$ with $\tw(G)\leq k-1$ and $\Delta(G)\leq d$,  
for any set $S\subseteq V(G)$ with $\alpha  k\leq|S| \leq 3 \alpha   kd$, there exists a tree-partition $(B_x:x\in V(T))$ of $G$ such that:
\begin{itemize}
    \item $|B_x|\leq \frac{3\alpha(\alpha-1)}{\alpha-2} kd - \frac{\alpha}{\alpha-2} k$ for each $x\in V(T)$,
    \item $\Delta(T)\leq \frac{3 \alpha}{\alpha-2}d + \frac{\alpha-4}{\alpha-2}$, 
    \item $|V(T)|\leq \tfrac{2}{\alpha k}|V(G)|$.
\end{itemize}
Moreover, there exists $z\in V(T)$ such that:
\begin{itemize}
    \item $S\subseteq B_z$, 
    \item $|B_z|\leq \frac{\alpha-1}{\alpha-2} |S|- \frac{\alpha}{\alpha-2} k$,
    \item $\deg_T(z)\leq \frac{1}{(\alpha-2)k} |S| - \frac{2}{\alpha-2}$.
\end{itemize}
\end{lem}

\begin{proof}
We proceed by induction on $|V(G)$|.



\textbf{Case 1.} $|V(G-S)|\leq 
\frac{3\alpha(\alpha-1)}{\alpha-2} kd - \frac{\alpha}{\alpha-2} k$: Let $T$ be the tree with $V(T)=\{y,z\}$ and $E(T)=\{yz\}$. Note that 
$|V(T)|=2\leq \frac{2}{\alpha k}|S| \leq \tfrac{2}{\alpha k}|V(G)|$ and $\Delta(T)=1\leq \frac{3 \alpha}{\alpha-2} d + \frac{\alpha-4}{\alpha-2}$ and $\deg_T(z)=1\leq \frac{1}{(\alpha-2)k}|S|-\frac{2}{\alpha-2}$. Let $B_z:=S$ and $B_y:=V(G-S)$. Thus $|B_z|=|S|\leq \frac{\alpha-1}{\alpha-2} |S|- \frac{\alpha}{\alpha-2} k\leq \frac{3\alpha(\alpha-1)}{\alpha-2} kd - \frac{\alpha}{\alpha-2} k$ and $|B_y|\leq |V(G-S)|\leq \frac{3\alpha(\alpha-1)}{\alpha-2}     kd$. Hence $(B_x:x\in V(T))$ is the desired tree-partition of $G$. 

Now assume that $|V(G-S)|\geq 
\frac{3\alpha(\alpha-1)}{\alpha-2} kd - \frac{\alpha}{\alpha-2} k$.

\textbf{Case 2.} $\alpha k \leq |S|\leq 3 \alpha   k$: Let $S':=\bigcup_{v\in S}N_G(v)\setminus S$. Thus $|S'|\leq d |S|\leq 3 \alpha   kd$. If $|S'|< \alpha  k$ then add $\alpha k-|S'|$ vertices from $V(G-S-S')$ to $S'$, so that $|S'|=\alpha k$. This is well-defined since 
$|V(G-S)| \geq 
\frac{3\alpha(\alpha-1)}{\alpha-2} kd - \frac{\alpha}{\alpha-2} k \geq \alpha k$, implying $|V(G-S-S')| \geq \alpha k-|S'|$.
By induction, there exists a tree-partition $(B_x:x\in V(T'))$ of $G-S$ such that:
\begin{itemize}
    \item $|B_x|\leq \frac{3\alpha(\alpha-1)}{\alpha-2} kd - \frac{\alpha}{\alpha-2} k$ for each $x\in V(T')$,
    \item $\Delta(T')\leq \frac{3 \alpha}{\alpha-2}d + \frac{\alpha-4}{\alpha-2}$, 
    \item $|V(T')|\leq \frac{2}{\alpha k}|V(G-S)|$.
\end{itemize}
Moreover, there exists $z'\in V(T')$ such that:
\begin{itemize}
    \item $S'\subseteq B_{z'}$, 
    \item $|B_{z'}|\leq \frac{\alpha-1}{\alpha-2} |S'|- \frac{\alpha}{\alpha-2} k \leq \frac{3 \alpha (\alpha-1)}{\alpha-2}   kd - \frac{\alpha}{\alpha-2} k$,
    \item $\deg_{T'}(z')\leq \frac{1}{(\alpha-2)k} |S'| - \frac{2}{\alpha-2} \leq  \frac{3 \alpha}{\alpha-2}  d -  \frac{2}{\alpha-2} $.
\end{itemize}
Let $T$ be the tree obtained from $T'$ by adding one new node $z$ adjacent to $z'$. Let $B_z:=S$. So $(B_x:x\in V(T))$ is a tree-partition of $G$ with width at most 
\[\max\{
\tfrac{3\alpha(\alpha-1)}{\alpha-2} kd - \tfrac{\alpha}{\alpha-2} k,|S|\} \leq \max\{
\tfrac{3\alpha(\alpha-1)}{\alpha-2} kd - \tfrac{\alpha}{\alpha-2} k,3 \alpha  k\} = 
\tfrac{3\alpha(\alpha-1)}{\alpha-2} kd - \tfrac{\alpha}{\alpha-2} k.\] By construction, 
\begin{align*}
\deg_T(z) & = 1 \leq \tfrac{1}{(\alpha-2)k} |S| -  \tfrac{2}{\alpha-2} \quad \text{and }\\
\deg_{T}(z')  & = \deg_{T'}(z')+1 \leq \tfrac{3 \alpha}{\alpha-2}   d -  \tfrac{2}{\alpha-2}  + 1  = \tfrac{3 \alpha}{\alpha-2}    d
+ \tfrac{\alpha-4}{\alpha-2}. 
\end{align*}
Every other vertex in $T$ has the same degree as in $T'$. Hence $\Delta(T)\leq \frac{3 \alpha}{\alpha-2}d + \frac{\alpha-4}{\alpha-2}$, as desired. 
Also $|V(T)|=|V(T')|+1\leq \frac{2}{\alpha k}|V(G-S)| +1 < \tfrac{2}{\alpha k}|V(G)|$ since $|S|\geq \alpha k$. 
Finally, $S=B_z$ and $|B_z|=|S| \leq \frac{\alpha-1}{\alpha-2}  |S|- \frac{\alpha}{\alpha-2} k$.

\textbf{Case 3.} $3 \alpha   k \leq |S|\leq 3 \alpha  kd$: By the separator lemma of \citet[(2.6)]{RS-II}, there are induced subgraphs $G_1$ and $G_2$ of $G$ with $G_1\cup G_2=G$ and $|V(G_1\cap G_2)|\leq k$, where $|S\cap V(G_i)|\leq \frac23 |S|$ for each $i\in\{1,2\}$. Let $S_i := (S\cap V(G_i))\cup V(G_1\cap G_2)$ for each $i\in\{1,2\}$.

We now bound $|S_i|$. For a lower bound, since $|S\cap V(G_1)|\leq \frac23 |S|$, we have $|S_2|\geq |S\setminus V(G_1)|\geq \frac13 |S| \geq \frac13 3 \alpha  k \geq \alpha k $. By symmetry, $|S_1|\geq  \alpha k $. For an upper bound, $|S_i|\leq\frac23 |S| + k \leq 2 \alpha  kd + k \leq 3 \alpha  kd$. Also note that $|S_1|+|S_2| \leq |S|+2|V(G_1\cap G_2)| \leq |S|+2k$. We have shown that $\alpha k \leq |S_i|\leq 3 \alpha  kd$ for each $i\in\{1,2\}$. Thus we may apply induction to $G_i$ with $S_i$ the specified set. Hence there exists a tree-partition $(B^i_x:x\in V(T_i))$ of $G_i$ such that:
\begin{itemize}
    \item $|B^i_x|\leq \frac{3\alpha(\alpha-1)}{\alpha-2} kd - \frac{\alpha}{\alpha-2} k$ for each $x\in V(T_i)$,
    \item $\Delta(T_i)\leq \frac{3 \alpha}{\alpha-2}d + \frac{\alpha-4}{\alpha-2} \leq \frac{|V(G_i)|}{2k}$, 
    \item $|V(T_i)|\leq \frac{2}{\alpha k}|V(G_i)|$.
\end{itemize}
Moreover, there exists $z_i\in V(T_i)$ such that:
\begin{itemize}
    \item $S_i\subseteq B_{z_i}$, 
    \item $|B_{z_i}|\leq \frac{\alpha-1}{\alpha-2} |S_i|- \frac{\alpha}{\alpha-2} k$,
    \item $\deg_{T_i}(z_i)\leq \frac{1}{(\alpha-2)k} |S_i| -  \frac{2}{\alpha-2} $.
\end{itemize}
Let $T$ be the tree obtained from the disjoint union of $T_1$ and $T_2$ by identifying $z_1$ and $z_2$ into a vertex $z$. Let $B_z:= B^1_{z_1}\cup B^2_{z_2}$. Let $B_x:= B^i_x$ for each $x\in V(T_i)\setminus\{z_i\}$. Since $G=G_1\cup G_2$ and $V(G_1\cap G_2) = B^1_{z_1}\cap B^2_{z_2} \subseteq B_z$, we have that $(B_x:x\in V(T))$ is a tree-partition of $G$. By construction, $S\subseteq B_z$ and since $V(G_1\cap G_2)\subseteq B^i_{z_i}$ for each $i$, 
\begin{align*}
    |B_z| 
    & \leq |B^1_{z_1}|+|B^2_{z_2}| - |V(G_1\cap G_2)|\\
    & \leq (\tfrac{\alpha-1}{\alpha-2} |S_1|- \tfrac{\alpha}{\alpha-2} k) +  (\tfrac{\alpha-1}{\alpha-2} |S_2|- \tfrac{\alpha}{\alpha-2} k) - |V(G_1\cap G_2)|\\
    & = \tfrac{\alpha-1}{\alpha-2} ( |S_1|+ |S_2|) - \tfrac{2\alpha}{\alpha-2} k - |V(G_1\cap G_2)|\\
    & \leq \tfrac{\alpha-1}{\alpha-2} ( |S| + 2|V(G_1\cap G_2)| ) - \tfrac{2\alpha}{\alpha-2} k - |V(G_1\cap G_2)|\\
    & = \tfrac{\alpha-1}{\alpha-2}  |S|  - \tfrac{2\alpha}{\alpha-2} k + \tfrac{\alpha}{\alpha-2}  |V(G_1\cap G_2)|\\
    & \leq \tfrac{\alpha-1}{\alpha-2}  |S|  - \tfrac{2\alpha}{\alpha-2} k + \tfrac{\alpha}{\alpha-2}  k\\
    & = \tfrac{\alpha-1}{\alpha-2} |S| -  \tfrac{\alpha}{\alpha-2}  k \\
    & \leq \tfrac{3\alpha(\alpha-1)}{\alpha-2}     kd -  \tfrac{\alpha}{\alpha-2}  k .
\end{align*}
Every other part has the same size as in the tree-partition of $G_1$ or $G_2$. So this tree-partition of $G$ has width at most $\frac{3\alpha(\alpha-1)}{\alpha-2} kd - \frac{\alpha}{\alpha-2} k$.  Note that 
\begin{align*}
 \deg_T(z)  & = \deg_{T_1}(z_1) + \deg_{T_2}(z_2)\\
    & \leq  \tfrac{1}{(\alpha-2)k}|S_1|- \tfrac{2}{\alpha-2} + \tfrac{1}{(\alpha-2)k}|S_2|- \tfrac{2}{\alpha-2} \\
    & \leq  \tfrac{1}{(\alpha-2)k}(|S_1|+|S_2|)- \tfrac{4}{\alpha-2} \\
    & \leq  \tfrac{1}{(\alpha-2)k}(|S|+2k) -
    \tfrac{4}{\alpha-2} \\
    & \leq  \tfrac{1}{(\alpha-2)k}|S|- \tfrac{2}{\alpha-2} \\
    & \leq  \tfrac{3 \alpha}{\alpha-2} d- \tfrac{2}{\alpha-2} \\
    & < \tfrac{3 \alpha}{\alpha-2}     d + \tfrac{\alpha-4}{\alpha-2}. 
\end{align*}
Every other node of $T$ has the same degree as in $T_1$ or $T_2$. 
Thus $\Delta(T) \leq \tfrac{3 \alpha}{\alpha-2}    d + \tfrac{\alpha-4}{\alpha-2}$. Finally, 
\begin{align*}
    |V(T)|  =|V(T_1)|+|V(T_2)|-1 
    & \leq \tfrac{2}{\alpha k}|V(G_1)| + \tfrac{2}{\alpha k}|V(G_2)|-1\\
    & \leq  \tfrac{2}{\alpha k}(|V(G)|+k)-1\\
    & < \tfrac{2}{\alpha k}|V(G)|.
\end{align*}
This completes the proof.
\end{proof}

A proof directly analogous to the proof of \cref{ImprovedTreeProduct} using 
\cref{alpha} with $\alpha=4$ instead of \cref{heart} implies the following slight strengthening of \cref{ImprovedTreeProduct}.

\begin{thm}
For $k\in\NN$, every non-trivial graph $G$ with $\tw(G)\leq k-1$ has a $T$-partition of width at most \[2\tw(G)\,(9\Delta(G)-1),\] 
for some tree $T$ with $\Delta(T)\leq 6\Delta(G)$ and $|V(T)|\leq \max\{\frac{|V(G)|}{2k},1\}$.
\end{thm}

Similarly, \cref{alpha} with $\alpha=2+\sqrt{2}$ (chosen to minimise $\frac{3\alpha(\alpha-1)}{\alpha-2}$) leads to the next result.

\begin{thm}
\label{ImprovedTreeProductSqrt}
For $k\in\NN$, every non-trivial graph $G$ with $\tw(G)\leq k-1$ has a $T$-partition of width at 
\[(1+\sqrt{2})k\,(3(1+\sqrt{2})\Delta(G) -1),\]
for some tree $T$ with $\Delta(T)\leq (3+3\sqrt{2})\Delta(G)-3(\sqrt{2}-1)$ and $|V(T)|\leq \max\{\tfrac{2|V(G)|}{(2+\sqrt{2}) k},1\}$. 
\end{thm}

{\fontsize{11pt}{12pt}\selectfont
\bibliographystyle{DavidNatbibStyle}
\bibliography{DavidBibliography}}
\end{document}